# Chirped amplitude mode in photo-excited superconductors

### v2

@@ II Model and method
We solve the problem within time-dependent dynamical mean-field theory (DMFT) on the Bethe lattice with infinite coordination number, using the Nambu formalism to describe the s-wave superconducting phase. This leads to matrix-valued Green’s functions $G_{\alpha\beta}(t,t^{\prime})=-i\langle\mathcal{T}\psi_{\alpha}(t)\psi^{\dagger}_{\beta}(t^{\prime})\rangle,$ with the Nambu spinor $\psi=\{d_{\uparrow},d^{\dagger}_{\downarrow}\}$. For the impurity solver, we use the self-consistent second-order perturbation theory known as second Born, or GF2 [48, 49, 50]. We refer to Appendix A for detailed implementation information. Collective order parameter dynamics take place on long time scales, and such long-time integration of the KBE is made possible by using a compressed representation of the Green’s function to reduce both the computational and memory complexity [42].
@@ III Results
To excite the amplitude mode, we pump the system with a pulse which has a frequency twice the band gap ($t_{c}^{\mathrm{Pump}}=16$, $\sigma=6.5$, $\omega=4\Delta$). The time evolution of the order parameter is presented in Fig. 1. At weak excitation strength, the order parameter decays at an extremely slow exponential rate [51, 52]. The expected thermal values $\phi_{th}$, marked by the horizontal lines, are not reached on the time scale of our simulation; see also Fig. 1(c), in which the thermal values are subtracted. For intermediate pump strengths, we observe the decay of the order parameter through several orders of magnitude. On top of the exponential relaxation, well-defined oscillations correspond to the amplitude mode excitation [53, 10]. As the pump strength is increased, the initial amplitude of these oscillations increases; however, they also have shorter lifetimes. Increasing the pump amplitude further, we reach a point at which the lifetimes of these oscillations hinder our ability to study their frequency [51, 54, 55], see Fig. 6(a,b) in Appendix B.
@@
Whereas the dissipationless BCS dynamics lead to either power-law decay [25, 56] or persistent oscillations [27, 29, 30, 28] of the superconducting amplitude depending on the initial condition, the inclusion of fluctuations modifies the Higgs mode dynamics into an exponential decay, see Appendix B. This is consistent with the effect of thermal noise on the dissipationless BCS dynamics [27, 34]. The lifetime of the amplitude mode decays with increasing pump strength (see Fig. 2(a-d)), and changes by more than an order of magnitude from the weak to the strong excitation regime.
The chirping of the amplitude mode is the main result of this work. Observing such phenomena requires microscopic treatment of fluctuations beyond the mean-field theory, see Appendix B for comparison, as the resulting state is highly nonthermal. Moreover, the chirping phenomenon is not limited to superconductors but should be observed in many systems with broken symmetry. For example, the superconducting solution analyzed in this work can be exactly mapped to the evolution of the antiferromagnetic state in the repulsive Hubbard model [57], and it is degenerated with the charge density waves. Both of these phases have been addressed experimentally  [58, 59, 60, 61]. In actual experimental setups, there is a trade-off in pump strength between the chirping velocity and the lifetime of the amplitude mode, which depends on the details of the pulse and the material studied.
@@
Based on these results, we propose that the direct measurement of the photoinduced current is a convenient way to measure both the chirping of the amplitude mode as well as the decay envelope of the Higgs mode which would be power-law (exponential) in the dissipationless (fluctuating) description. The photo-induced current and chirping can be measured in the pump-probe setup using either free-space electro-optic sampling, like in Ref. [11], or transmission line experiments, which were recently used to detect superconducting nonlinear transport [67], the light-induced anomalous Hall effect due to circularly polarized light [68], and ultrafast resistive switching in 1T-TaS2 [69]. In these experiments, one measures the variation of the transmitted field induced by the probe field in the material. The modification of the transmitted field originates from the variation of the microscopic polarization in the material which is connected with the microscopic currents[70, 71, 72].
@@
Since our analysis was enabled by the HODLR compression technique for the numerical solution of the KBE [42], we consider the compressibility of the dynamics of symmetry-broken states, and the performance of the scheme compared with the direct solution of KBE. In Fig. 4(a), we observe approximately quadratic scaling of the computational cost with the propagation time over a range of pump intensities and equilibrium state temperatures. We reach 32768 time steps in roughly one day on a single-core of a workstation using 0.5 GB of memory. By contrast, we extrapolate that cubic-scaling direct time stepping using the NESSi code [35] would have taken 129 days with 137 GB of memory. Fig. 4(b) shows the different scaling between the full Green’s function and the HODLR representation. At $N=32768$ we have a factor of $50$-$200$ memory compression of the Green’s function and self-energy compared with direct storage on a two-time grid. Appendix C considers the compressibility in more detail.
@@ Appendix A Implementation details
We use the fifth-order backward differentiation formula multistep method to discretize the KBE. Gregory quadrature is used to discretize memory integrals to fifth-order accuracy. The discrete Lehmann representation [83], implemented in libdlr [84], is used to discretize the imaginary time variables appearing in the vertical leg of the Keldysh contour, as in Ref. [85]. We have confirmed that the compression of the Green’s function using the HODLR scheme preserves conservation laws, namely the density and energy, to within the Singular Value Decomposition truncation tolerance used in the scheme, which we set to $10^{-8}$.

### v3

@@ II Model and method
We solve the problem within time-dependent dynamical mean-field theory (DMFT) on the Bethe lattice with infinite coordination number, using the Nambu formalism to describe the s-wave superconducting phase. This leads to matrix-valued Green’s functions $G_{\alpha\beta}(t,t^{\prime})=-i\langle\mathcal{T}\psi_{\alpha}(t)\psi^{\dagger}_{\beta}(t^{\prime})\rangle,$ with the Nambu spinor $\psi=\{d_{\uparrow},d^{\dagger}_{\downarrow}\}$. For the impurity solver, we use the self-consistent second-order perturbation theory known as second Born, or GF2 [48, 49, 50]. Collective order parameter dynamics take place on long time scales, and such long-time integration of the KBE is made possible by using a compressed representation of the Green’s function to reduce both the computational and memory complexity [42]. We refer to the SM for further implementation details.
@@ III Results
To excite the amplitude mode, we pump the system with a pulse which has a frequency twice the band gap ($t_{c}^{\mathrm{Pump}}=16$, $\sigma=6.5$, $\omega=4\Delta$). The time evolution of the order parameter is presented in Fig. 1. At weak excitation strength, the order parameter decays at an extremely slow exponential rate [51, 52]. The expected thermal values $\phi_{th}$, marked by the horizontal lines, are not reached on the time scale of our simulation; see also Fig. 1(c), in which the thermal values are subtracted. For intermediate pump strengths, we observe the decay of the order parameter through several orders of magnitude. On top of the exponential relaxation, well-defined oscillations correspond to the amplitude mode excitation [53, 10]. As the pump strength is increased, the initial amplitude of these oscillations increases; however, they also have shorter lifetimes. Increasing the pump amplitude further, we reach a point at which the lifetimes of these oscillations hinder our ability to study their frequency [51, 54, 55], see Fig. 6(a,b) in the SM.
@@
Whereas the dissipationless BCS dynamics lead to either power-law decay [25, 56] or persistent oscillations [27, 29, 30, 28] of the superconducting amplitude depending on the initial condition, the inclusion of fluctuations modifies the Higgs mode dynamics into an exponential decay, see SM. This is consistent with the effect of thermal noise on the dissipationless BCS dynamics [27, 57]. The lifetime of the amplitude mode decays with increasing pump strength (see Fig. 2(a-d)), and changes by more than an order of magnitude from the weak to the strong excitation regime.
The chirping of the amplitude mode is the main result of this work. Observing such phenomena requires microscopic treatment of fluctuations beyond the mean-field theory, see SM for comparison, as the resulting state is highly nonthermal. Moreover, the chirping phenomenon is not limited to superconductors but should be observed in many systems with broken symmetry. For example, the superconducting solution analyzed in this work can be exactly mapped to the evolution of the antiferromagnetic state in the repulsive Hubbard model [58], and it is degenerated with the charge density waves. Both of these phases have been addressed experimentally  [59, 60, 61, 62]. In actual experimental setups, there is a trade-off in pump strength between the chirping velocity and the lifetime of the amplitude mode, which depends on the details of the pulse and the material studied.
@@
Based on these results, we propose that the direct measurement of the photo-induced current is a convenient way to measure both the chirping of the amplitude mode as well as to decay envelope of the Higgs mode which would be power-law (exponential) in the dissipationless (fluctuating) description. The photo-induced current and chirping can be measured in the pump-probe setup using either free-space electro-optic sampling, like in Ref. [11], or transmission line experiments, which were recently used to detect superconducting nonlinear transport [68], the light-induced anomalous Hall effect due to circularly polarized light [69], and ultrafast resistive switching in 1T-TaS2 [70]. In these experiments, one measures the variation of the transmitted field induced by the probe field in the material. The modification of the transmitted field originates from the variation of the microscopic polarization in the material which is connected with the microscopic currents[71, 72, 73].
@@
Since our analysis was enabled by the HODLR compression technique for the numerical solution of the KBE [42], we consider the compressibility of the dynamics of symmetry-broken states, and the performance of the scheme compared with the direct solution of KBE. In Fig. 4(a), we observe approximately quadratic scaling of the computational cost with the propagation time over a range of pump intensities and equilibrium state temperatures. We reach 32768 time steps in roughly one day on a single-core of a workstation using 0.5 GB of memory. By contrast, we extrapolate that cubic-scaling direct time stepping using the NESSi code [35] would have taken 129 days with 137 GB of memory. Fig. 4(b) shows the different scaling between the full Green’s function and the HODLR representation. At $N=32768$ we have a factor of $50$-$200$ memory compression of the Green’s function and self-energy compared with direct storage on a two-time grid. The SM considers the compressibility in more detail.
@@ Appendix A Implementation details
We use the fifth-order backward differentiation formula multistep method to discretize the KBE. Gregory quadrature is used to discretize memory integrals to fifth-order accuracy. The discrete Lehmann representation [84], implemented in libdlr [85], is used to discretize the imaginary time variables appearing in the vertical leg of the Keldysh contour, as in Ref. [86]. We have confirmed that the compression of the Green’s function using the HODLR scheme preserves conservation laws, namely the density and energy, to within the SVD truncation tolerance used in the scheme, which we set to $10^{-8}$.